# DIAS: A Comprehensive Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

### 12 Dec 2023 v2

@@ III-B2 Scribble Annotation
Scribble-based image segmentation has proven to be an efficacious and sought-after method for WSS [45]. This approach to weak annotation allows annotators to easily mark object centers without the challenging task of delineating class boundaries. To explore the potential of scribble supervision for IA segmentation in DSA sequences, we generated two forms of scribble annotations for the training set of DIAS: the Scribbles Annotation with Little Clinical Experience (SALE) and Random Drawing based the Full Annotation (RDFA). SALE simulates vessel annotation with little expert experience, wherein the thick vessels highlighted by the contrast agent are annotated, as illustrated by the red marker in Fig. 4 (a). Despite receiving clinicians’ basic training, accurately distinguishing the thin vessels and edges of thick blood vessels with low contrast remains challenging. Consequently, pixels in these regions were annotated during annotation. Moreover, scribble segmentation relies on supervised information for every category. Apart from annotating the vessels, we conducted extensive scribble annotations for the background wherever possible, as illustrated by the blue marker in Fig. 4 (a). Follow [46], we generate RDFA by skeletonizing and randomly pruning each category of the full annotation, as shown in Fig. 4 (b). RDFA simulates random scribble annotations, diverging from SALE, which focuses on easier annotations. In SALE and RDFA, we set the pixel values of the background and vessel annotations to 0 and 1, respectively, while pixels without annotations are assigned a value of 255.
@@ IV Methods
Automatic IA segmentation in DSA sequences is an under-explored and challenging task. To fully explore potential research topics, three elaborately designed benchmarks for IA segmentation were proposed, aiming to provide comprehensive benchmarks for FSS, WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
@@ IV-A Fully-supervised Segmentation
Segmenting IA in DSA sequence can be considered as a dimensionality reduction (DR) task. Given an image $I\in\mathbb{R}^{\prod_{d=1}^{N}n_{d}}$, where $N$ represents the number of dimensions and $n_{d}$ is the size of the image in the corresponding dimension $d$, DR segmentation can be defined as finding a function $f$ such that $f:I\rightarrow O$, where $O\in\mathbb{R}^{\prod_{d=1}^{M}n_{d}}$ represents segmentation output and $M$ ($<N$) is its dimensionality [47]. For IA segmentation in DSA sequence, $N=3$, $M=2$. Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network (VSS-Net) efficiently captures spatio-temporal representation of vessels from various scales to achieve precise IA segmentation in DSA. Figure 5 (a,b) illustrates an overview of the proposed network architecture. VSS-Net is composed of multi-scale sequence feature extraction modules (SFEM). These modules handle the 2D+T input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations. Subsequently, these feature maps can be fed into any 2D model following the encoder-decoder framework, enabling vessel segmentation in the sequence.
@@
In the encoder, the input feature map from each stage is integrated with the raw sequence, serving to supplement fine-grained semantic information, except for the initial stage. Specifically, the raw sequence has been resized to be consistent with the corresponding feature maps by bilinear interpolation. Each image of sequence are concatenated with feature maps from the same temporal sequence, and then they are fused using a 1x1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels. This process is formulated as:
$Q=\sum_{i=1}^{N}W\ast(x_{i}\oplus f_{i})$ (1)
@@
where, $x$ represents the input frame, $f$ is the feature map, $N$ denotes the length of the sequence. $W$ are $1\times 1$ convolutional kernels. $\ast$ denotes the convolution operation. $\oplus$ denotes the concatenation operation.
@@
Subsequently, we employ a bidirectional ConvGRU to construct the spatiotemporal connection between the feature maps of sequence. ConvGRU [48] is an extension of the traditional GRU that integrates convolutional operations, making it more suitable for spatial data like images. Consider $x_{t}$ as the input to the ConvGRU, with corresponding hidden states denoted by $h_{t}$, we can delineate the updating process of ConvGRU unrolled by time as follows:
$z_{t}$ $=\sigma(W_{z}\ast x_{z}+U_{z}\ast h_{t-1})$ (2)
$r_{t}$ $=\sigma(W_{r}\ast x_{r}+U_{r}\ast h_{t-1})$ (3)
$\tilde{h}_{t}$ $=\tanh(W_{h}\ast x_{t}+U_{h}\ast(r_{t}\odot h_{t-1}))$ (4)
$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (5)
@@
where, $\sigma$ represents the sigmoid function. $\tanh$ represents the hyperbolic tangent function. $\odot$ denotes the Hadamard product, i.e., element-wise multiplication. $W_{r}$, $W_{z}$, $W_{h}$ and $U_{r}$, $U_{z}$, $U_{h}$ are 2D-convolutional kernels.
@@
$h_{t}^{f}$ $=ConvGRU(h_{t-1},x_{t})$ (6)
$h_{t}^{b}$ $=ConvGRU(h_{t+1},x_{t})$ (7)
$h_{t}$ $=\tanh(W_{f}\ast h_{t}^{f}+W_{b}\ast h_{t}^{b})$ (8)
where, $h_{t}^{f}$ and $h_{t}^{b}$ denote the hidden states from the forward and backward ConvGRU units, respectively. $x_{t}$ is the $t^{th}$ output feature from WSC.
@@ IV-B Weakly-supervised Segmentation
A common strategy for scribble-supervised semantic segmentation is to handle the problem as a fully supervised task, using a loss function $\mathcal{L}$ on the available labeled points. However, this baseline approach does not utilize the unlabeled points that contain critical semantic information. The performance degradation is unavoidable due to the lack of labeled data points. To enhance the segmentation performance of scribble, the key lies in extending the supervision of labeled pixels to unlabeled pixels. To overcome these issues, we propose SSCR, a novel framework for learning-based scribble image segmentation composed of scribble supervision and consistency regularization (CR). Fig. 5 (b) presents the implementation framework for SSCR. SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing segmentation network using mutual consistency and cross pseudo supervision (CPS).
@@
$\mathcal{L}_{pce}(y,\hat{y})=-\sum_{i=1}^{M}\sum_{j=1}^{N}{\mathbbm{1}(p_{i,j}\in S)(y_{i,j}\log(\hat{y}_{i,j|\theta})})$ (9)
where, $M$ and $N$ represent the width and height of the image, respectively. $\hat{y}_{i,j|\theta}$ is the predict for the pixel $p_{i,j}\in S$ given the network parameters $\theta$, and $y_{i,j}$ denotes the ground truth label.
@@
where, $s_{i,j}\in S$ defines annotated pixels, and $\mathcal{H}(p_{i,j};\theta)$ defines the predicted output of $p_{i,j}$. $C$ is the channel of annotated class, i.e., 0 and 1.
@@
$\mathcal{L}_{scr}=$ $\frac{1}{2}(\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL))$ (11)
$\mathcal{L}_{cps}=$ $\frac{1}{2}(\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A}))$ (12)
$\mathcal{L}_{con}=$ $\mathcal{L}_{mse}(PL_{A},PL_{B})$ (13)
where, $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function. $P_{A}$ and $P_{B}$ are the outputs of two models. $PL_{A}$ and $PL_{B}$ are the pseudo labels. $SL$ is the scribble label.
Finally, the training objective $\mathcal{L}$ is formulated as
$\mathcal{L}=\mathcal{L}_{scr}+\lambda_{1}\mathcal{L}_{cps}+\lambda_{2}\mathcal{L}_{con}$ (14)
where, $\lambda_{1}$ and $\lambda_{2}$ are the weight of loss components $\mathcal{L}_{cps}$ and $\mathcal{L}_{con}$.
@@ IV-C Semi-supervised Segmentation
Annotation of cerebral vessels manually from DSA image sequences is a laborious and time-consuming task. Semi-supervised learning has shown the potential to alleviate this problem by learning from a large set of unlabeled images and limited labeled samples [42]. The primary challenge in semi-supervised setting lies in how to effectively utilize the unlabeled images. Self-training [50] via pseudo labeling is a conventional, simple, and popular pipeline to leverage unlabeled data, where the retrained student is supervised with hard labels produced by the teacher trained on labeled data, which is commonly regarded as a form of entropy minimization in semi-supervised learning. For semi-supervised segmentation problem, let $\mathcal{D}^{l}=\{(x_{i},y_{i})\}_{i=1}^{M}$ be pixel-wise labeled images and $\mathcal{D}^{u}=\{u_{i}\}_{i=1}^{N}$ be unlabeled images, where in most cases $N\gg M$. Self-training can be simplified into the following steps:
  1. [leftmargin=*,itemsep=-5pt]
@@
We follow the plainest self-training strategy to provide a SSS benchmark for DIAS. The injection of strong data augmentations on unlabeled images is highly advantageous as it decouples the predictions as well as alleviates overfitting on noisy pseudo-labels. Despite its simplicity, self-training with strong data augmentations surpasses existing methods without any bells and whistles [41]. We performed two types of data augmentations in our implementation: weak and strong, denoted by $A^{s}$ and $A^{w}$ respectively. Weak data augmentations (WDA) consist of horizontal flipping, vertical flipping, and $[90,180,270]$ degree rotation. Strong data augmentations (SDA) include Cutout, adding white Gaussian noise, Gaussian blurring, and elastic deformation. We randomly adjust the contrast of each image in the DSA sequence to simulate angiography with intensity differences. We randomly extract patches from the DSA sequence to train the teacher and student models. Notably, the patches in each mini-batch were randomly extracted from $\mathcal{D}^{l}$ and $\mathcal{\hat{D}}^{u}$ with a probability $P$ when retraining the student model. Additionally, the final performance can be further improved by switching the teacher-student role and relabeling unlabeled images for iterative training. The pseudocode of the random patch-based self-training framework, referred to as RPST, is provided in Algorithm 1.
@@
          Weak/strong augmentation $\mathcal{A}^{w}/{A}^{s}$,
@@ V-A Implementation Details and Evaluation Metrics
The DIAS dataset was divided into 30 training samples, 10 validation samples, and 20 testing samples on the patient level. To facilitate input into the model, all DSA sequences were resampled to a length of 8. All DSA sequences were z-score normalized. Patches of size $64\times 64$ were randomly cropped for training, with a batch size of 64. The model parameters were optimized using AdamW [51] with an initial learning rate of 5e-4. The learning rate is gradually reduced by the cosine annealing algorithm over 100 epochs. The parameter $P$ was set to 0.5, i.e., half of the patches in each batch were extracted from labeled data, while the rest came from unlabeled data in SSS experiment. For SSCR, the weights of loss components $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 1 and 0.5 . We only use one network to generate the results for evaluation. All experiments used PyTorch with constant hyperparameters and were conducted on a single GeForce RTX 3090 GPU. We calculated the dice similarity coefficient (DSC), the area under the receiver operating characteristic curve (AUC), accuracy (Acc), sensitivity (Sen), specificity (Spe), and intersection over union (IOU) to evaluate the performance of DSA sequence segmentation. Moreover, we introduced the vascular connectivity (VC) [12] metric, which calculates the ratio of connected components in the segmentation image to those in the ground truth, as a means of measuring the connectivity of vessels.
V-B Fully-supervised Segmentation
[…]
$y=\mathcal{H}_{2d}(Squ(x))$ (15)
where, $Squ$ is dimensionality squeeze operation and $\mathcal{H}_{2d}$ is a 2D model.
$y=Conv_{1\times 1}(Squ(\mathcal{H}_{3d}(x)))$ (16)
Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC and VC, among all the evaluated methods. Using FR-UNet as a baseline, VSS-Net shows improvements in numerous metrics except for Spe. Specifically, Sen has increased by 1.3%. ST U-Net, also employing ConvGRU, shows performance very close to VSS-Net, achieving the best results in Acc and Spe. However, the number of model parameters is approximately 5.58 times that of VSS-Net. The VC score that is 19.57 higher than VSS-Net indicates poorer vascular connectivity. Both IPN and SVS-Net employ 3D convolutions in the encoder and 2D convolutions in the decoder, incorporating dimensionality reduction in the intermediate stages. PSC’s segmentation results are similar to those of 3D models, but IPN shows poorer performance, possibly due to the absence of skip connections, which are significantly important for enhancing the detailing in vascular structures. Upon overall comparison of the DSC scores for all ‘DR+2D model’ and ‘DR+3D model’, 2D models slightly outperform 3D models with fewer parameters. Although in terms of data format, 2D+time DSA sequences are equivalent to 3D data, they do not have a one-to-one spatial relationship with the ground truth. More importantly, 3D convolutions can only model local temporal representations. The ‘DR+2D model’ benefits from treating ‘sequences’ as ‘channels’, which enables it to learn global representations along the sequence dimension while performing 2D convolutions, which aggregates contrast agent responses that reflect vascular morphological information in the sequence direction. However, the approach of setting fixed channel numbers is not applicable to variable image sequences. We have visualized the segmentation results of the methods that exhibited outstanding performance, as showed in Figure 6. In the figure, the red and blue colors represent pixels that are false negatives and false positives, respectively, indicative of segmentation errors. It can be observed that a prevalent issue shared by all methods is the abundant incorrect segmentation of thin vessels. Specifically, we enlarged representative details in the images, as illustrated by the yellow and blue boxes in Figure 6. In comparison with other methods, VSS-Net has the highest number of white pixels in the fine vessels, indicating the best segmentation results.
@@ V-B Fully-supervised Segmentation
2D UNet [27] 31.04 0.7677 0.9647 0.757 0.9829 0.6258 0.9811 68.29
Res-UNet [54] 9.37 0.7656 0.9655 0.7425 0.985 0.6235 0.9816 60.56
UNet++ [53] 9.05 0.7768 0.9659 0.7726 0.983 0.6374 0.983 54.65
@@
VSS-Net 18.69 0.7822 0.9668 0.777 0.9838 0.6442 0.9857 43.02
Figure 6: Visualization of segmentation results of state-of-the-art models. The first row is coronal view, and the second row is sagittal view. The red pixel indicates false negatives, and the green pixel indicates false positives.

### 17 Dec 2023 v3

@@ III-B2 Scribble Annotation
Scribble-based image segmentation has proven to be an efficacious and sought-after method for WSS [45]. This approach to weak annotation allows annotators to easily mark object centers without the challenging task of delineating class boundaries. To explore the potential of scribble supervision for IA segmentation in DSA sequences, we generated two forms of scribble annotations for the training set of DIAS: the Scribbles Annotation with Little Clinical Experience (SALE) and Random Drawing based on the Full Annotation (RDFA). SALE simulates vessel annotation with little expert experience, wherein the thick vessels highlighted by the contrast agent are annotated, as illustrated by the red marker in Fig. 4 (a). Despite receiving clinicians’ basic training, accurately distinguishing the thin vessels and edges of thick blood vessels with low contrast remains challenging. Consequently, pixels in these regions were not annotated during annotation. Moreover, scribble segmentation relies on supervised information for every category. Apart from annotating the vessels, we conducted extensive scribble annotations for the background wherever possible, as illustrated by the blue marker in Fig. 4 (a). Follow [46], we generate RDFA by skeletonizing and randomly pruning each category of the full annotation, as shown in Fig. 4 (b). RDFA simulates random scribble annotations, diverging from SALE, which focuses on easier annotations. In SALE and RDFA, we set the pixel values of the background and vessel annotations to 0 and 1, respectively, while pixels without annotations are assigned a value of 255.
@@ IV Methods
Automatic IA segmentation in DSA sequences is an under-explored and challenging task. To fully explore potential research topics, an elaborately designed benchmark for IA segmentation was proposed, aiming to provide comprehensive evaluation of the DIAS dataset using FSS, WSS, and SSS, which are increasingly drawing attention in the medical image analysis community.
@@ IV-A Fully-supervised Segmentation
Segmenting IA in DSA sequence can be considered as a dimensionality reduction (DR) task. Given an image $I\in\mathbb{R}^{\prod_{d=1}^{M}n_{d}}$, where $M$ represents the number of dimensions and $n_{d}$ is the size of the image in the corresponding dimension $d$, DR segmentation can be defined as finding a function $f$ such that $f:I\rightarrow O$, where $O\in\mathbb{R}^{\prod_{d=1}^{N}n_{d}}$ represents segmentation output and $N$ ($<M$) is its dimensionality [47]. For IA segmentation in DSA sequence, $M=3$, $N=2$.
@@
Contrary to traditional medical image segmentation models that solely focus on spatial features, our proposed vessel sequence segmentation network (VSS-Net) efficiently captures spatio-temporal representations of vessels across various scales to achieve precise IA segmentation in DSA. Figure 5 (a,b) illustrates an overview of the proposed network architecture. VSS-Net comprises multi-scale sequence feature extraction modules (SFEM). These modules handle the 2D+Time input sequences and incorporate feature maps from the preceding stage, resulting in 2D feature maps that capture both spatial and temporal representations. Subsequently, these feature maps can be fed into any 2D model following the encoder-decoder framework, enabling vessel segmentation in the sequence.
@@
In the encoder, the input feature map from each stage is integrated with the raw sequence, serving to supplement fine-grained semantic information, except for the initial stage. Specifically, the raw sequence has been resized to be consistent with the corresponding feature maps by bilinear interpolation. Each image of sequence are concatenated with feature maps from the same temporal sequence, and then they are fused using a 1$\times$1 convolution with shared weights (WSC), leading to a reduction by half in the number of channels. This process is formulated as:
$Q=\sum_{i=1}^{N}W\ast(x_{i}\oplus f_{i}),$ (1)
@@
where $x$ represents the input frame, $f$ is the feature map, $N$ denotes the length of the sequence. $W$ are $1\times 1$ convolutional kernels. $\ast$ denotes the convolution operation. $\oplus$ denotes the concatenation operation.
@@
Subsequently, we employ a bidirectional ConvGRU to construct the spatiotemporal connection between the feature maps of the sequence. ConvGRU [48] is an extension of the traditional GRU that integrates convolutional operations, making it more suitable for spatial data like images. Consider $x_{t}$ as the input to the ConvGRU, with corresponding hidden states denoted by $h_{t}$, we can delineate the updating process of ConvGRU unrolled by time as follows:
$z_{t}$ $=\sigma(W_{z}\ast x_{z}+U_{z}\ast h_{t-1}),$ (2)
$r_{t}$ $=\sigma(W_{r}\ast x_{r}+U_{r}\ast h_{t-1}),$ (3)
$\tilde{h}_{t}$ $=\tanh(W_{h}\ast x_{t}+U_{h}\ast(r_{t}\odot h_{t-1})),$ (4)
$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t},$ (5)
@@
where $\sigma$ represents the sigmoid function. $\tanh$ represents the hyperbolic tangent function. $\odot$ denotes the Hadamard product, i.e., element-wise multiplication. $W_{r}$, $W_{z}$, $W_{h}$ and $U_{r}$, $U_{z}$, $U_{h}$ are 2D-convolutional kernels.
@@
$h_{t}^{f}$ $=ConvGRU(h_{t-1},x_{t}),$ (6)
$h_{t}^{b}$ $=ConvGRU(h_{t+1},x_{t}),$ (7)
$h_{t}$ $=\tanh(W_{f}\ast h_{t}^{f}+W_{b}\ast h_{t}^{b}),$ (8)
where $h_{t}^{f}$ and $h_{t}^{b}$ denote the hidden states from the forward and backward ConvGRU units, respectively. $x_{t}$ is the $t^{th}$ output feature from WSC.
@@ IV-B Weakly-supervised Segmentation
A prevalent strategy for scribble-supervised semantic segmentation is to handle the problem as a fully supervised task, using a loss function $\mathcal{L}$ on the available labeled pixels. However, this baseline approach fails to harness the unlabeled points that contain critical semantic information. Consequently, the absence of sufficient labeled data invariably results in performance degradation. To enhance the segmentation performance of scribble, the key lies in extending the supervision of labeled pixels to unlabeled pixels. To overcome these issues, we propose SSCR, a novel framework for learning-based scribble image segmentation composed of scribble supervision and consistency regularization (CR). Fig. 5 (c) presents the implementation framework for SSCR. SSCR generates pseudo-annotations for the unlabeled pixels with guidance provided by scribble supervision, utilizing partial cross-entropy to extend supervised information, and maximizes the performance of scribble segmentation by regularizing segmentation network using mutual consistency and cross pseudo supervision (CPS).
@@
$\mathcal{L}_{pce}(y,\hat{y})=-\sum_{i=1}^{M}\sum_{j=1}^{N}{\mathbbm{1}(p_{i,j}\in S)(y_{i,j}\log(\hat{y}_{i,j|\theta})}),$ (9)
where $M$ and $N$ represent the width and height of the image, respectively. $\hat{y}_{i,j|\theta}$ is the predict for the pixel $p_{i,j}\in S$ given the network parameters $\theta$, and $y_{i,j}$ denotes the ground truth label.
@@
where $s_{i,j}\in S$ defines annotated pixels, and $\mathcal{H}(p_{i,j};\theta)$ defines the predicted output of $p_{i,j}$. $C$ is the channel of annotated class, i.e., 0 and 1.
@@
$\mathcal{L}_{scr}=$ $\frac{1}{2}(\mathcal{L}_{pce}(P_{A},SL)+\mathcal{L}_{pce}(P_{B},SL)),$ (11)
$\mathcal{L}_{cps}=$ $\frac{1}{2}(\mathcal{L}_{ce}(P_{A},PL_{B})+\mathcal{L}_{ce}(P_{B},PL_{A})),$ (12)
$\mathcal{L}_{con}=$ $\mathcal{L}_{mse}(PL_{A},PL_{B}),$ (13)
where $\mathcal{L}_{ce}$ denotes the cross-entropy loss function and $\mathcal{L}_{mse}$ denotes the mean square error loss function. $P_{A}$ and $P_{B}$ are the outputs of two models. $PL_{A}$ and $PL_{B}$ are the pseudo labels. $SL$ is the scribble label.
Finally, the training objective $\mathcal{L}$ is formulated as:
$\mathcal{L}=\mathcal{L}_{scr}+\lambda_{1}\mathcal{L}_{cps}+\lambda_{2}\mathcal{L}_{con},$ (14)
where $\lambda_{1}$ and $\lambda_{2}$ are the weight of loss components $\mathcal{L}_{cps}$ and $\mathcal{L}_{con}$.
@@ IV-C Semi-supervised Segmentation
Annotation of cerebral vessels manually from DSA image sequences is a laborious and time-consuming task. Semi-supervised learning has shown the potential to alleviate this problem by learning from a large set of unlabeled images and limited labeled samples [42]. The primary challenge in semi-supervised setting lies in how to effectively utilize the unlabeled images. Self-training [50] via pseudo labeling is a conventional, simple, and popular pipeline to leverage unlabeled data, where the retrained student is supervised with hard labels produced by the teacher trained on labeled data. For semi-supervised segmentation problem, let $\mathcal{D}^{l}=\{(x_{i},y_{i})\}_{i=1}^{M}$ be pixel-wise labeled images and $\mathcal{D}^{u}=\{u_{i}\}_{i=1}^{N}$ be unlabeled images, where in most cases $N\gg M$. Self-training can be simplified into the following steps:
  1. [leftmargin=*,itemsep=-4pt]
@@
As shown in Fig. 5 (d), we follow the plainest self-training strategy to provide a SSS benchmark for DIAS. The injection of strong data augmentations on unlabeled images is highly advantageous as it decouples the predictions as well as alleviates overfitting on noisy pseudo-labels. Despite its simplicity, self-training with strong data augmentations surpasses existing methods without any bells and whistles [41]. We performed two types of data augmentations in our implementation: weak and strong, denoted by $A^{w}$ and $A^{s}$ respectively. Weak data augmentations (WDA) consist of horizontal flipping, vertical flipping, and $[90,180,270]$ degree rotation. Strong data augmentations (SDA) include Cutout, adding white Gaussian noise, Gaussian blurring, and elastic deformation. Furthermore, we propose the application of random contrast adjustments to each image within the sequence to simulate angiography with intensity differences. We randomly extract patches from the DSA sequence to train the teacher and student models. Notably, the patches in each mini-batch were randomly extracted from $\mathcal{D}^{l}$ and $\mathcal{\hat{D}}^{u}$ with a probability $P$ when retraining the student model. Additionally, the final performance can be further improved by switching the teacher-student role and relabeling unlabeled images for iterative training. The pseudocode of the random patch-based self-training framework, referred to as RPST, is provided in Algorithm 1.
@@
          Weak/strong augmentation $\mathcal{A}^{w}/\mathcal{A}^{s}$,
@@ V-A Implementation Details and Evaluation Metrics
2D UNet [27] 31.04 0.7677 0.9647 0.7570 0.9829 0.6258 0.9811 68.29
Res-UNet [51] 9.37 0.7656 0.9655 0.7425 0.9850 0.6235 0.9816 60.56
UNet++ [52] 9.05 0.7768 0.9659 0.7726 0.9830 0.6374 0.9830 54.65
@@
VSS-Net 18.69 0.7822 0.9668 0.7770 0.9838 0.6442 0.9857 43.02
The DIAS dataset was divided into 30 training samples, 10 validation samples, and 20 testing samples on the patient level. To facilitate input into the model, all DSA sequences were resampled to a length of 8. Furthermore, z-score normalization was applied across all DSA sequences for standardization. Patches of size $64\times 64$ were randomly cropped for training, with a batch size of 64. The model parameters were optimized using AdamW [55] with an initial learning rate of 5e-4. The learning rate is gradually reduced by the cosine annealing algorithm over 100 epochs. The parameter $P$ was set to 0.5 in the RPST, i.e., half of the patches in each batch were extracted from labeled data, while the rest came from unlabeled data in SSS experiment. For the SSCR, the weights of loss components $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 1 and 0.5. We only use one network to generate the results for evaluation. All experiments used PyTorch with constant hyperparameters and were conducted on a single GeForce RTX 3090 GPU. We calculated the dice similarity coefficient (DSC), the area under the receiver operating characteristic curve (AUC), accuracy (Acc), sensitivity (Sen), specificity (Spe), and intersection over union (IOU) to evaluate the performance of DSA sequence segmentation. Moreover, we introduced the vascular connectivity (VC) [12] metric, which calculates the ratio of connected components in the segmentation image to those in the ground truth, as a means of measuring the connectivity of vessels.
V-B Fully-supervised Segmentation
[…]
$y=\mathcal{H}_{2d}(Squ(x)),$ (15)
where $Squ$ is dimensionality squeeze operation and $\mathcal{H}_{2d}$ is a 2D model.
$y=Conv_{1\times 1}(Squ(\mathcal{H}_{3d}(x))).$ (16)
Table II demonstrates that the comprehensive performance of VSS-Net surpasses other networks, achieving the best DSC, Sen, IOU, AUC and VC, among all the evaluated methods. Using FR-UNet as a baseline, VSS-Net shows improvements in numerous metrics except for Spe. Specifically, Sen has increased by 1.3%. ST U-Net, also employing ConvGRU, shows performance very close to VSS-Net, achieving the best results in Acc and Spe. However, the number of model parameters is approximately 5.58 times that of VSS-Net. The VC score that is 19.57 higher than VSS-Net indicates poorer vascular connectivity. Both IPN and SVS-Net employ 3D convolutions in the encoder and 2D convolutions in the decoder, incorporating dimensionality reduction in the intermediate stages. PSC’s segmentation results are similar to those of 3D models, but IPN shows poorer performance, possibly due to the absence of skip connections, which are significantly important for enhancing the detailing in vascular structures.
Figure 6: Visualization of segmentation results of state-of-the-art models. The first and third rows display images of the anteroposterior view, while the second and fourth rows present images of the lateral view. In the segmentation maps for various methods, red pixels indicate false negatives, and green pixels signify false positives. The white and black pixels represent correctly segmented vessels and background, respectively.
Upon overall comparison of the DSC scores for all ‘DR+2D model’ and ‘DR+3D model’, 2D models slightly outperform 3D models with fewer parameters. Although in terms of data format, 2D+Time DSA sequences are equivalent to 3D data, they do not have a one-to-one spatial relationship with the ground truth. More importantly, 3D convolutions can only model local temporal representations. The ‘DR+2D model’ benefits from treating ‘sequences’ as ‘channels’, which enables it to learn global representations along the sequence dimension while performing 2D convolutions, which aggregates contrast agent responses that reflect vascular morphological information in the sequence direction. However, the approach of setting fixed channel numbers is not applicable to variable image sequences. We have visualized the segmentation results of the methods that exhibited outstanding performance, as showed in Fig. 6. In the figure, the red and blue colors represent pixels that are false negatives and false positives, respectively, indicative of segmentation errors. It can be observed that a prevalent issue shared by all methods is the abundant incorrect segmentation of thin vessels. Specifically, we enlarged representative details in the images, as illustrated by the yellow and blue boxes in Fig. 6. In comparison with other methods, VSS-Net has the highest number of white pixels in the fine vessels, indicating the best segmentation results.